OpenStudy (anonymous):

how do you solve: (4e^(3x-1))= 24

OpenStudy (saifoo.khan):

\[e^{3x-1} = \frac{24}{4}\] \[e^{3x-1} = 6\] \[3x-1 = \ln 6\] im sure u can solve further.

OpenStudy (anonymous):

thank you. :D lol simply math sucks bleh XP heheheh

OpenStudy (saifoo.khan):

You welcome. xD